### v1

@@ 4 Subgroups of the Galois group, subfields, divisors and curvettes
Further on we shall discuss only Cases I and II. Let $(E_{\sigma}\circ E_{\delta})$, $\sigma,\delta\in\Gamma$, be the intersection matrix of the components $E_{\sigma}$ (infinite one in Case II). The diagonal entry $E_{\sigma}\circ E_{\sigma}$ is the self-intersection number of the component $E_{\sigma}$ on the final modification in Case I and on a modification at a step, after which the points of $E_{\sigma}$ are not blown-up, in Case II. One has $E_{\sigma}\circ E_{\sigma}<0$. The non-diagonal entry $E_{\sigma}\circ E_{\delta}$ is equal to 1 (respectively to 00) if these components intersect (do not intersect respectively).
Let $(m_{\sigma,\delta}):=-(E_{\sigma}\circ E_{\delta})^{-1}$. In Case II one can compute the entry $m_{\sigma,\delta}$ at a sufficiently late step (after which it stabilizes). In other words one can compute the intersection matrix and the inverse of it at each step and use their stabilization.
The number $m_{\sigma,\delta}$ can be interpreted as the intersection number of some germs of curves on $(\mathbb{C}^{2},0)$, so-called curvettes. A curvette correponding to the component $E_{\sigma}$ is the blow-down $C_{\sigma}$ of a complex-analytic smooth curve $\gamma_{\sigma}$ (in Case II on a modification where $E_{\sigma}$ stabilizes) transversal to the component $E_{\sigma}$ at a smooth point of the exceptional divisor, that is not an intersection point with other components. One can show that $m_{\sigma,\delta}$ is equal to the intersection number $C_{\sigma}\circ C_{\delta}$.
For $\sigma\in\Gamma$, let $m_{\sigma}$ be the value of $\nu_{C}$ on the left-hand side of an equation $f_{\sigma}=0$ defining the curvette $C_{\sigma}$, i. e., the intersection number $C\circ C_{\sigma}$. One can see that, in Case I, $m_{\sigma}$ is equal to $m_{\sigma,\delta_{C}}$ for $\delta_{C}$ such that, on the surface of the resolution, the strict transform of the branch $C$ intersects the component $E_{\delta_{C}}$ and, in Case II, $m_{\sigma}$ is equal to $m_{\sigma,\delta}$ for $\delta$ on the subgraph of the usual resolution of the curve $C$ far enough from the initial component of the modification.
@@
Let $G_{\sigma}$ be the isotropy subgroup of the component $E_{\sigma}$, i. e., the isotropy subgroup of the vertex $\sigma$ in $\Gamma$. The subgroup $G_{\sigma}$ has a finite index in the Galois group $G$. Let $\mathbb{K}_{\sigma}\subset\mathbb{C}$ be the invariant subfield of the group $G_{\sigma}$. One can see that, for $g\in G$, $G_{g\sigma}=gG_{\sigma}g^{-1}$, $\mathbb{K}_{g\sigma}=g\mathbb{K}_{\sigma}$. For all the vertices $\sigma\in\Gamma$ inbetween the initial one $\sigma_{0}=$1 and $\rho_{1}$ ($\rho_{1}$ included) the subgroups $G_{\sigma}$ and the subfields $\mathbb{K}_{\sigma}$ are the same: $G_{\sigma}=G=G_{0}$, $\mathbb{K}_{\sigma}=\mathbb{K}=\mathbb{K}_{0}$. For all vertices $\sigma\in\Gamma$ inbetween the vertex $\rho_{j}$ and $\rho_{j+1}$ ($\rho_{j}$ excluded, $\rho_{j+1}$ included) the isotropy subgroups $G_{\sigma}$ and the corresponding subfields $\mathbb{K}_{\sigma}$ are the same: $G_{\sigma}=G_{j}$, $\mathbb{K}_{\sigma}=\mathbb{K}_{j}$. In Case I, for all the vertices $\sigma\in\Gamma$ betwin $\rho_{s}$ and $\delta_{C}$ one has $G_{\sigma}=G_{s}$, $\mathbb{K}_{\sigma}=\mathbb{K}_{s}$. For vertices from $G$-orbits of the discussed ones, the corresponding subgroups are conjugate to $G_{\sigma}$ and the subfields are obtained from $\mathbb{K}_{\sigma}$ by the shift by an element of the Galois group $G$. One has $G=G_{0}\supset G_{1}\supset\ldots\supset G_{s}$, $\mathbb{K}=\mathbb{K}_{0}\subset\mathbb{K}_{1}\subset\ldots\subset\mathbb{K}_{s}$. One has $[K_{j}:K_{j-1}]=[G_{j-1}:G_{j}]$.
@@
On all the components $E_{\sigma}$ of the exceptional divisor except $E_{\rho_{j}}$ and those from its G-orbit, on each step the blow-up is made at one point with the coordinates $(0,a)$ in a standard chart with $a\in\mathbb{K}_{\sigma}$. In this case it can be convenient to change the coordinate $w$ by $\widetilde{w}=w/(w-a)$ or $\widetilde{w}=(w-a)/w$ so that the coordinates of the intersection points with the components on the geodesic between $\sigma_{0}$ and $\delta_{C}$ become equal to 00 and $\infty$. On the component $E_{\rho_{j}}$, the blow-ups are made at a certain point $(0,a)$ with $a\in\mathbb{K}_{j}\setminus\mathbb{K}_{j-1}$ (that is not from $\mathbb{K}_{\rho_{j}}=\mathbb{K}_{j-1}$) and at the points from its $G_{j-1}$-orbit. In this case a change as above is inappropriate.
@@
In what follows, we shall use curvettes of special type. In a standard chart covering an affine part of $E_{\sigma}$, let the (smooth and transversal) curve $\gamma_{\sigma}$ be defined over the field $\mathbb{K}_{\sigma}$. In other words, the curve $\gamma_{\sigma}$ is invariant under the action of $G_{\sigma}$. In particular one can take the curve $\{u=\tau$, $w=c=\text{const}\}$ in a standart chart with $c\in\mathbb{K}_{\sigma}$ ($u=0$ is the equation of $E_{\sigma}$ in the chart).
@@ 5 The semigroup Poincaré series of a curve valuation
Let $\nu(=\nu_{C})$ be the valuation on ${\mathcal{E}}_{\mathbb{K}^{2},0}$ defined by a curve $C\subset(\mathbb{C}^{2},0)$ and assume that $\nu_{C}$ is a valuation of the type I or II described in Section 3. Let $P^{S}_{C}(t)=\sum_{m\in S^{\mathbb{K}}_{\nu}}t^{m}$ be the semigroup Poincaré series of $\nu_{C}$ (see Definition 2).
For $\sigma\in\check{\Gamma}$, let $M_{\sigma}$ be the value of $\nu$ on the left-hand side of the equation $F_{\sigma}=0$ definig a $G_{\sigma}$-curvette at the component $E_{\sigma}$. One can see that $M_{\sigma}=\sum_{[g]\in G/G_{\sigma}}m_{g\sigma}$. (Compare with the definition of $M_{\sigma}$ in [7], where $\mathbb{K}=\mathbb{R}$.)
@@
Let $S_{C}$ be the usual semigroup of the complex curve $C$, i. e., $S_{C}=\{\nu(f):f\in{\mathcal{O}}_{\mathbb{C}^{2},0}\}$. The set of multiplicities $\{m_{\sigma_{0}},\ldots,m_{\sigma_{g}}\}$ is the minimal set of generators of $S_{C}$. Let $e_{i}:=\gcd\{m_{\sigma_{0}},\,\ldots,m_{\sigma_{i}}\}$ for $i=0,1,\ldots,g$ and let $N_{i}=e_{i-1}/e_{i}$ for $i=1,2,\ldots,g$. (One has $m_{\tau_{i}}=N_{i}m_{\sigma_{i}}$ for $i=1,2,\ldots,g$: see, e. g., [12] or [2].)
@@
For $j=1,\ldots,s$, let $\ell_{j}:=[\mathbb{K}_{j}:\mathbb{K}_{j-1}]=[G_{j-1}:G_{j}]:=|\frac{G_{j-1}}{G_{j}}|$.
@@
Let $\omega\in\check{\Gamma}$ be such that $\rho_{d}<\omega\leq\rho_{d+1}$ and $\tau_{p}\leq\omega<\tau_{{p}+1}$. One has
@@
for some non negative integers $k_{i}$ $(0\leq i\leq{p})$ and positive integers $r_{j}$ $(1\leq j\leq d)$.
@@
Moreover, it is known that $m_{\omega}\in\langle m_{0},\,\ldots,m_{p}\rangle$ and so $m_{\omega}=k_{0}m_{0}+\sum_{i=1}^{p}k_{i}m_{i}$, for some $k_{0}\geq 0$ and $0\leq k_{i}<N_{i}$ for $i\geq 1$ (which are unique with this restrictions). Then one can write
@@
Let us assume that $q<{p}$. For $j\leq d$ one has $e_{\rho_{j}}(\varphi_{\omega})/e_{\tau_{{p}-1}}(\varphi_{\omega})=e_{\rho_{j}}(\varphi_{{p}})/e_{\tau_{{p}-1}}(\varphi_{p})$, moreover $e_{\tau_{{p}-1}}(\varphi_{p})=1$, so $e_{\rho_{j}}(\varphi_{\omega})=e_{\tau_{{p}-1}}(\varphi_{\omega})e_{\rho_{j}}(\varphi_{p})$. It is known that $e_{i}=e_{\tau_{i}}(\varphi_{C})$, $i\leq g$ and $\varphi_{C}$ is an equation of the complex curve $C$. As a consequence, for $i\leq p$, one has $N_{i}=e_{i-1}/e_{i}=e_{\tau_{i-1}}(\varphi_{\omega})/e_{\tau_{i}}(\varphi_{\omega})$ and so
$e_{\rho_{j}}(\varphi_{\omega})=N_{p}e_{\tau_{p}}(\varphi_{\omega})e_{\rho_{j}}(\varphi_{p})\;.$
The same computations applied to $\varphi_{i}$ (instead of $\varphi_{\omega}$) and $\varphi_{i-1}$ (for $\tau_{i}>\tau_{i-1}>\rho_{j}$) show that $e_{\rho_{j}}(\varphi_{i})=N_{i-1}e_{\tau_{i-1}}(\varphi_{i})e_{\rho_{j}}(\varphi_{i-1})=N_{i-1}e_{\rho_{j}}(\varphi_{i-1})$.
Using this equality in the rigth part of the equation (6) and also the expression of $M_{i}$ above, one gets:
@@
with $k^{\prime}_{p}=N_{p}e_{\tau_{p}}(\varphi_{\omega})-k_{p}\geq 1$, and so
@@
In the same way, using repeatedly that $e_{\rho_{i}}(\varphi_{j})=N_{i-1}e_{\rho_{j}}(\varphi_{i-1})$ for $i\geq q$ one gets:
$M_{\omega}=(\sum_{0}^{q-1}k_{i}m_{i})+\sum_{i=q}^{p}k_{i}M_{i}+r_{d}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}$ (7)
for some integer $r_{d}\geq 1$. Now, using that $L^{d}_{j}=[G_{d-1}:G_{d}]L^{d-1}_{j}$ for $j\leq d-1$ in the last summand of the above equality (7), one has
$r_{d}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}=r_{d}[G_{d-1}:G_{d}]\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}+r_{d}L^{d}_{d}e_{\rho_{d}}(\varphi_{q})m_{\rho_{d}}\;.$
Using that $L^{d}_{d}=[G_{d-1}:G_{d}]-1$, $e_{\rho_{j}}(\varphi_{q})=e_{\rho_{j}}(\varphi_{\rho_{d}})$, $e_{\rho_{d}}(\varphi_{\rho_{d}})=1$ and the corresponding expression for $M_{\rho_{d}}$ in the same terms as at the beginning of the proof for $M_{\omega}$, the last equality can be written as:
@@
Thus, taking into account that $\ell_{j}=[G_{j-1}:G_{j}]$ and the last equality, the equation (7) for $M_{\omega}$ becomes:
@@
If $d=1$, one has $m_{i}=M_{i}$ for $i=0,\ldots,q-1$ and the proof is finished. Otherwise, we apply the induction hypothesis to the last two summands in (8) (in brackets). Notice that the factor $r_{d}\geq 1$ does not affects the proof above and the terms $e_{\rho_{j}}(\varphi_{q})$ (corresponding to a curvette at $E_{\sigma_{q}}$) could be treated (arithmetically) as we did with $e_{\rho_{j}}(\varphi_{\omega})$ for $j<d$. As a consequence, we have:
@@
For $\sigma\in\check{\Gamma}$, $\tau_{k}\leq\sigma<\tau_{k+1}$ it is known that $m_{\sigma}\in\langle m_{\sigma_{0}},\ldots,m_{\sigma_{k}}\rangle$ or equivalently that $e_{k}$ divides $m_{\sigma}$ (see [12]. Taken into account the expression for $M_{i}$ given in equation (5), this implies that $e_{i-1}$ divides $\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{i}}$ and so $\gcd(M_{0},\,\ldots,M_{i})=\gcd(m_{0},\,\ldots,m_{i})=e_{i}$ and
@@
Lemma 2.
For $i=1,\ldots,g$ one has $(N_{i}-1)M_{\sigma_{i}}\notin\langle M_{\sigma_{0}},\ldots,M_{\sigma_{i-1}}\rangle$, $N_{i}M_{\sigma_{i}}\in\langle M_{\sigma_{0}},\ldots,M_{\sigma_{i-1}}\rangle$ and $N_{i-1}M_{\sigma_{i-1}}<M_{\sigma_{i}}$. In particular, $\{M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\}$ is the minimal set of generators of $S=\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ and, moreover, $S$ is the semigroup of values of a complex analytic irreducible curve.
@@
Note that, as a consequence of the Remark 5, $(N_{i}-1)M_{i}\notin\langle M_{0},\ldots,M_{i-1}\rangle$.
Let us prove that for $i=1,\ldots,g$, one has that $N_{i-1}M_{i-1}<M_{i}$. If $\rho_{d}\leq\tau_{i-1}<\tau_{i}$, the equation (5) for $M_{i}$ together with the equality $e_{\rho_{i}}=N_{i-1}e_{\rho_{i}}(\varphi_{i-1})$ for $i\leq d$ implie that
@@
$L^{d}_{j}=[G_{d-1}:G_{d}]L^{d-1}_{j}=\ell_{d-1}L^{d-1}_{j}=\ell_{d-1}\cdots\ell_{d^{\prime}}L^{d^{\prime}}_{j}$

### v2

@@ 4 Subgroups of the Galois group, subfields, divisors and curvettes
Further on we shall discuss only Cases I and II. We do not discuss Case III since in this case the valuation $\nu_{C}$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$ is a multiple of a divisorial valuation whose Poincaré series is described in Section 7. Let $(E_{\sigma}\circ E_{\delta})$, $\sigma,\delta\in\Gamma$, be the intersection matrix of the components $E_{\sigma}$ (infinite one in Case II). The diagonal entry $E_{\sigma}\circ E_{\sigma}$ is the self-intersection number of the component $E_{\sigma}$ on the final modification in Case I and, in Case II, a modification at a step, after which the points of $E_{\sigma}$ are not blown-up. One has $E_{\sigma}\circ E_{\sigma}<0$. The non-diagonal entry $E_{\sigma}\circ E_{\delta}$ is equal to 1 (respectively to 0) if these components intersect (do not intersect respectively).
Let $(m_{\sigma,\delta}):=-(E_{\sigma}\circ E_{\delta})^{-1}$. In order not to treat infinite matrices, in Case II the entry $m_{\sigma,\delta}$ can be computed at a step of the resolution process after which the components $E_{\sigma}$ and $E_{\delta}$ are not modified.
The number $m_{\sigma,\delta}$ can be interpreted as the intersection number of some germs of curves on $(\mathbb{C}^{2},0)$, so-called curvettes. A curvette corresponding to the component $E_{\sigma}$ is the blow-down $C_{\sigma}$ of a complex-analytic smooth curve $\gamma_{\sigma}$ (in Case II on a modification after which the component $E_{\sigma}$ is not modified) transversal to the component $E_{\sigma}$ at a smooth point of the exceptional divisor, that is not an intersection point with other components. One can show that $m_{\sigma,\delta}$ is equal to the intersection number $C_{\sigma}\circ C_{\delta}$.
For $\sigma\in\Gamma$, let $m_{\sigma}=\nu_{C}(f_{\sigma})$, where $f_{\sigma}=0$ is an equation defining a curvette $C_{\sigma}$. It is equal to the intersection number $C\circ C_{\sigma}$. One can see that, in Case I, $m_{\sigma}$ is equal to $m_{\sigma,\delta_{C}}$ for $\delta_{C}$ such that, on the surface of the resolution, the strict transform of the branch $C$ intersects the component $E_{\delta_{C}}$. In Case II, the vertex $\delta_{C}$ makes sense only at any step of the resolution process of the curve $C$. One can see that $m_{\sigma}=m_{\sigma,\delta}$ at a step of the resolution process when the branch $C$ itself is resolved (not in the $G$-invariant sense) and after which the component $E_{\sigma}$ is not modified.
@@
Let, as above, $G_{\sigma}$ be the isotropy subgroup of the component $E_{\sigma}$, i. e., the isotropy subgroup of the vertex $\sigma$ in $\Gamma$. The subgroup $G_{\sigma}$ has a finite index in the Galois group $G$. Let $\mathbb{K}_{\sigma}\subset\mathbb{C}$ be the invariant subfield of the group $G_{\sigma}$. One can see that, for $g\in G$, $G_{g\sigma}=gG_{\sigma}g^{-1}$, $\mathbb{K}_{g\sigma}=g\mathbb{K}_{\sigma}$. For all the vertices $\sigma\in\Gamma$ in between the initial one $\sigma_{0}=$1 and $\rho_{1}$ ($\rho_{1}$ included) the subgroups $G_{\sigma}$ and the subfields $\mathbb{K}_{\sigma}$ are the same: $G_{\sigma}=G=G_{0}$, $\mathbb{K}_{\sigma}=\mathbb{K}=\mathbb{K}_{0}$. For all vertices $\sigma\in\Gamma$ in between the vertex $\rho_{j}$ and $\rho_{j+1}$ ($\rho_{j}$ excluded, $\rho_{j+1}$ included) the isotropy subgroups $G_{\sigma}$ and the corresponding subfields $\mathbb{K}_{\sigma}$ are the same: $G_{\sigma}=G_{j}$, $\mathbb{K}_{\sigma}=\mathbb{K}_{j}$. In Case I, for all the vertices $\sigma\in\Gamma$ between $\rho_{s}$ and $\delta_{C}$ one has $G_{\sigma}=G_{s}$, $\mathbb{K}_{\sigma}=\mathbb{K}_{s}$. For vertices from $G$-orbits of the discussed ones, the corresponding subgroups are conjugate to $G_{\sigma}$ and the subfields are obtained from $\mathbb{K}_{\sigma}$ by the shift by an element of the Galois group $G$. One has $G=G_{0}\supset G_{1}\supset\ldots\supset G_{s}$, $\mathbb{K}=\mathbb{K}_{0}\subset\mathbb{K}_{1}\subset\ldots\subset\mathbb{K}_{s}$. One has $\mathbb{K}_{j}:\mathbb{K}_{j-1}]=[G_{j-1}:G_{j}]$.
@@
On all the components $E_{\sigma}$ of the exceptional divisor except $E_{\rho_{j}}$ and those from its $G$-orbit, on each step the blow-up is made at one point with the coordinates $(0,a)$ in a standard chart with $a\in\mathbb{K}_{\sigma}$. In this case it can be convenient to change the coordinate $w$ by $\widetilde{w}=w/(w-a)$ or $\widetilde{w}=(w-a)/w$ so that the coordinates of the intersection points with the components on the geodesic between $\sigma_{0}$ and $\delta_{C}$ become equal to 0 and $\infty$. On the component $E_{\rho_{j}}$, the blow-ups are made at a certain point $(0,a)$ with $a\in\mathbb{K}_{j}\setminus\mathbb{K}_{j-1}$ (that is not from $\mathbb{K}_{\rho_{j}}=\mathbb{K}_{j-1}$) and at the points from its $G_{j-1}$-orbit. In this case a coordinate change as above cannot be used since it is not defined over $\mathbb{K}_{j-1}$.
@@
In what follows, we shall use curvettes of special type. In a standard chart covering an affine part of $E_{\sigma}$, let the (smooth and transversal) curve $\gamma_{\sigma}$ be defined over the field $\mathbb{K}_{\sigma}$. In other words, the curve $\gamma_{\sigma}$ is invariant under the action of $G_{\sigma}$. In particular one can take the curve $\{u=\tau$, $w=c=\text{const}\}$ in a standard chart with $c\in\mathbb{K}_{\sigma}$ ($u=0$ is the equation of $E_{\sigma}$ in the chart).
@@ 5 The semigroup Poincaré series of a curve valuation
Let $\nu(=\nu_{C})$ be the valuation on ${\mathcal{E}}_{\mathbb{K}^{2},0}$ defined by an irreducible curve $C\subset(\mathbb{C}^{2},0)$ and assume that $\nu_{C}$ is a valuation of the type I or II described in Section 3. Let $P^{S}_{C}(t)=\sum_{m\in S^{\mathbb{K}}_{\nu}}t^{m}$ be the semigroup Poincaré series of $\nu_{C}$ (see Definition 2).
For $\sigma\in\check{\Gamma}$, let $M_{\sigma}=\nu(F_{\sigma})$ where $F_{\sigma}=0$ is the equation defining a $G_{\sigma}$-curvette at the component $E_{\sigma}$. One can see that $M_{\sigma}=\sum_{[g]\in G/G_{\sigma}}m_{g\sigma}$. (Compare with the definition of $M_{\sigma}$ in [7], where $\mathbb{K}=\mathbb{R}$.)
@@
Let $S^{\mathbb{C}}_{C}$ be the usual semigroup of the complex branch $C$, i. e., $S^{\mathbb{C}}_{C}=\{\nu(f):f\in{\mathcal{O}}_{\mathbb{C}^{2},0}\}$. The set of multiplicities $\{m_{\sigma_{0}},\ldots,m_{\sigma_{g}}\}$ is the minimal set of generators of $S^{\mathbb{C}}_{C}$. Let $e_{i}:=\gcd\{m_{\sigma_{0}},\,\ldots,m_{\sigma_{i}}\}$ for $i=0,1,\ldots,g$ and let $N_{i}=e_{i-1}/e_{i}$ for $i=1,2,\ldots,g$. (One has $m_{\tau_{i}}=N_{i}m_{\sigma_{i}}$ for $i=1,2,\ldots,g$: see, e. g., [15] or [2].)
@@
For $j=1,\ldots,s$, let $\ell_{j}:=[\mathbb{K}_{j}:\mathbb{K}_{j-1}]=[G_{j-1}:G_{j}]=\#(G_{j-1}/{G_{j}})$.
@@
Let $\omega\in\check{\Gamma}$ be such that $\rho_{d}<\omega\leq\rho_{d+1}$ and $\tau_{p}<\omega\leq\tau_{{p}+1}$. One has
@@
for some non negative integers $k_{i}$ $(0\leq i\leq{p})$, $k_{i}<N_{i}$ for $i\geq 1$, and positive integers $r_{j}$ $(1\leq j\leq d)$.
@@
Moreover, it is known that $m_{\omega}\in\langle m_{0},\,\ldots,m_{p}\rangle$, where $\langle m_{0},\,\ldots,m_{p}\rangle$ stands for the subsemigroup of $\mathbb{Z}_{\geq 0}$ generated by $m_{0},\,\ldots,m_{p}$, i.e.
$\langle m_{0},\,\ldots,m_{p}\rangle:=\{n\in\mathbb{N}:n=a_{0}m_{0}+\cdots+a_{p}m_{p}\text{ with }a_{i}\in\mathbb{Z}_{\geq 0}\}\;,$
and so $m_{\omega}=k_{0}m_{0}+\sum_{i=1}^{p}k_{i}m_{i}$, for some $k_{0}\geq 0$ and $0\leq k_{i}<N_{i}$ for $i\geq 1$ (which are unique with these restrictions). Then one can write
@@
Let us assume that q<pq<{p}, note that in this case we have $\tau_{q-1}<\rho_{d}\leq\tau_{q}<\tau_{p}\leq\omega$ (see Figure 6). For $j\leq d$ one has $e_{\rho_{j}}(\varphi_{\omega})/e_{\tau_{{p}-1}}(\varphi_{\omega})=e_{\rho_{j}}(\varphi_{{p}})/e_{\tau_{{p}-1}}(\varphi_{p})$, moreover $e_{\tau_{{p}-1}}(\varphi_{p})=1$, so $e_{\rho_{j}}(\varphi_{\omega})=e_{\tau_{{p}-1}}(\varphi_{\omega})e_{\rho_{j}}(\varphi_{p})$. It is known that $e_{i}=e_{\tau_{i}}(\varphi_{C})$, $i\leq g$, where $\varphi_{C}=0$ is an equation of the complex curve $C$. As a consequence, for $i\leq p$, one has $N_{i}=e_{i-1}/e_{i}=e_{\tau_{i-1}}(\varphi_{\omega})/e_{\tau_{i}}(\varphi_{\omega})$ and so
$e_{\rho_{j}}(\varphi_{\omega})=N_{p}e_{\tau_{p}}(\varphi_{\omega})e_{\rho_{j}}(\varphi_{p})\;.$ (7)
For $i,j$ such that $\tau_{i}>\tau_{i-1}>\rho_{j}$, the same computations applied to $\varphi_{i}$ (instead of $\varphi_{\omega}$) and $\varphi_{i-1}$ (instead of $\varphi_{p}$) show that
$e_{\rho_{j}}(\varphi_{i})=N_{i-1}e_{\tau_{i-1}}(\varphi_{i})e_{\rho_{j}}(\varphi_{i-1})=N_{i-1}e_{\rho_{j}}(\varphi_{i-1})\;.$ (8)
Using equality (7) in the right part of equation (6), one gets:
@@
with $k^{\prime}_{p}=N_{p}e_{\tau_{p}}(\varphi_{\omega})-k_{p}\geq N_{p}-k_{p}\geq 1$, with $k_{p}$ being the integer appearing in equation (6). Now, using the expression of $M_{i}$ in (5) for $i=p$, one has
@@
In the same way, using repeatedly that $e_{\rho_{j}}(\varphi_{i})=N_{i-1}e_{\rho_{j}}(\varphi_{i-1})$ and the equation (5) for $M_{i}$, $i\geq q$, one gets:
$M_{\omega}=\sum_{i=0}^{q-1}k_{i}m_{i}+\sum_{i=q}^{p}k_{i}M_{i}+r_{d}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}$ (9)
for some integer $r_{d}\geq 1$. Now, one can use the equality $L^{d}_{j}=[G_{d-1}:G_{d}]L^{d-1}_{j}$ for $j\leq d-1$ and (8) in the last summand of (9) to get:
$r_{d}\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}=r_{d}[G_{d-1}:G_{d}]\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}+r_{d}L^{d}_{d}e_{\rho_{d}}(\varphi_{q})m_{\rho_{d}}\;.$ (10)
Using that $L^{d}_{d}=[G_{d-1}:G_{d}]-1$, $e_{\rho_{j}}(\varphi_{q})=e_{\rho_{j}}(\varphi_{\rho_{d}})$, $e_{\rho_{d}}(\varphi_{\rho_{d}})=1$ and the corresponding expression for $M_{\rho_{d}}$ in the same terms as we use at the beginning of the proof for $M_{\omega}$, the equality (10) can be written as:
@@
Thus, using this equality and taking into account that $\ell_{j}=[G_{j-1}:G_{j}]$, the equation (9) for $M_{\omega}$ becomes:
@@
If $d=1$, one has $m_{i}=M_{i}$ for $i=0,\ldots,q-1$ and the proof is finished. Otherwise, we repeat the above computations for the last two summands in (11) (in brackets):
$\sum_{0}^{q-1}k_{i}m_{i}+r_{d}\sum_{j=1}^{d-1}L^{d-1}_{j}e_{\rho_{j}}(\varphi_{q})m_{\rho_{j}}\;.$ (12)
To do that, note that the equation (5) remains valid for $M_{i}$, $i\leq q$, and that the equation (12) is essentially the same as the one in the right hand side of (6):
  1. 1.
    The factor $r_{d}\geq 1$ does not affect the computations in the proof.
  2. 2.
    Equation (8) could be applied to $\varphi_{p}$ (and to the succesive $\varphi_{i}$, $i<p$) instead of equation (7) for $\varphi_{\omega}$.
  3. 3.
    The succesive integers $k^{\prime}_{i}\geq N_{i}-k_{i}$ are always greater or equal to 1.
As a consequence, we get:
@@
Proposition 1.
The set of integers $\{M_{\sigma_{0}},\,\ldots,M_{\sigma_{g}}\}$ is a set of generators of the semigroup of values $S^{\mathbb{K}}_{C}$ of the valuation $\nu$ on ${\mathcal{E}}_{\mathbb{K}^{2},0}$.
Proof..
It is clear that the semigroup $S^{\mathbb{K}}_{\nu}$ is generated by the integers $M_{\sigma}$, for all $\sigma\in\check{\Gamma}$. Thus one has to show that $M_{\sigma}\in\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ for any $\sigma$.
Let $\alpha,\beta]$ denote the geodesic in $\check{\Gamma}$ between the vertices $\alpha$ and $\beta$. If $\omega\in[\sigma_{0},\delta_{C}]$, Lemma 1 gives
$M_{\omega}=\sum_{i=0}^{h}k_{i}M_{\sigma_{i}}+\sum_{j=1}^{d}r_{j}(\ell_{j}-1)M_{\rho_{j}}\;,$
where $\rho_{d}<\omega\leq\rho_{d+1}$, $\tau_{h}<\omega\leq\tau_{h+1}$, $k_{i}$ are non-negative integers and $r_{j}$ are positive integers. The same equation applied to $\rho_{j}$ gives us that $M_{\rho_{j}}\in\langle M_{0},\,\ldots,M_{i}\rangle$ if $\tau_{i}\leq\rho_{j}<\tau_{i+1}$. In particular, $M_{\rho_{j}}\in\langle M_{0},\,\ldots,M_{h}\rangle\subset\langle M_{0},\,\ldots,M_{g}\rangle$
@@
For $\sigma\in\check{\Gamma}$, $\tau_{k}<\sigma\leq\tau_{k+1}$, it is known that $m_{\sigma}\in\langle m_{\sigma_{0}},\ldots,m_{\sigma_{k}}\rangle$ or equivalently that $e_{k}$ divides $m_{\sigma}$ (see [15]). Taken into account the expression for $M_{i}$ given in equation (5), this implies that $e_{i-1}$ divides $\sum_{j=1}^{d}L^{d}_{j}e_{\rho_{j}}(\varphi_{i})m_{\rho_{i}}$ and so $\gcd(M_{0},\,\ldots,M_{i})=\gcd(m_{0},\,\ldots,m_{i})=e_{i}$ and
@@
Proposition 2.
For $i=1,\ldots,g$, one has $(N_{i}-1)M_{\sigma_{i}}\notin\langle M_{\sigma_{0}},\ldots,M_{\sigma_{i-1}}\rangle$, $N_{i}M_{\sigma_{i}}\in\langle M_{\sigma_{0}},\ldots,M_{\sigma_{i-1}}\rangle$ and $N_{i-1}M_{\sigma_{i-1}}<M_{\sigma_{i}}$. In particular, $\{M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\}$ is the minimal set of generators of the semigroup $S^{\mathbb{K}}_{C}:=\langle M_{\sigma_{0}},\ldots,M_{\sigma_{g}}\rangle$ and, moreover, $S^{\mathbb{K}}_{C}$ is the semigroup of values of a complex analytic irreducible curve.
@@
If $(N_{i}-1)M_{i}\in\langle M_{0},\,\ldots,M_{i-1}\rangle$ then, as a consequence of Remark 7, $e_{i-1}$ divides $(N_{i}-1)M_{i}$ and so $e_{i-1}$ divides $(N_{i}-1)m_{i}$. But this contradicts the properties of the minimal set of generators of the semigroup $S^{\mathbb{C}}_{C}$.
Note that $N_{i}M_{i}=M_{\tau_{i}}$, then $N_{i}M_{i}\in\langle M_{0},\,\ldots,M_{i-1}\rangle$ as a consequence of Proposition 1 (more precisely of the first part of its proof).
Let us prove that, for $i=1,\ldots,g$, one has that $N_{i-1}M_{i-1}<M_{i}$. If $\rho_{d}\leq\tau_{i-1}<\tau_{i}$, the equation (5) for $M_{i}$ together with the equality $e_{\rho_{i}}=N_{i-1}e_{\rho_{i}}(\varphi_{i-1})$ for $i\leq d$ imply that
@@
$L^{d}_{j}=[G_{d-1}:G_{d}]L^{d-1}_{j}=\ell_{d-1}L^{d-1}_{j}=\ell_{d-1}\cdots\ell_{d^{\prime}}L^{d^{\prime}}_{j}\;,$